OpenStudy (anonymous):

NEED ANSWER ASAP ! Create your own radical equation. Describe in complete sentences and demonstrate the process in finding its solution(s) .

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you mean any radical equation?

OpenStudy (anonymous):

this is the easiest one √4

OpenStudy (anonymous):

square root of 4

OpenStudy (anonymous):

√2√2 =√4

OpenStudy (anonymous):

√2 * √2 = √4

OpenStudy (anonymous):

is that good?

OpenStudy (anonymous):

why is that because there is some relation between x^2 since you are actually squaring the number the number you are squaring is √x^

OpenStudy (anonymous):

forgot its this √x^2

OpenStudy (anonymous):

is this good?

OpenStudy (anonymous):

well there are other cases where it is not just so simple as √x^2

OpenStudy (anonymous):

for example what would the square root of √28 be?

OpenStudy (anonymous):

√28 ^ √4 *√7 square root of 28 = square root 4 times 7 ^ √2 * 2 square root 4 = square root 2 times 2

OpenStudy (anonymous):

x^2 =x √x^2 <-this will be referred as the perfect square of a number if above statement is true

OpenStudy (anonymous):

okies i hope this helped